26th January 2006 Hoyle committee room 10.00 - 12:40 Present: DWE, MJI, PSB, STH, MR, JRL (RGM after 11:30) Apologies: NAW, EGS, WJS Agenda ------ 1. Actions from last meeting 2. Comments on WFAU minutes 3. VDFS VISTA update 4. WFCAM processing 5. Upcoming VDMT meeting and this quarter WPs 6. Data archives update 7. Optical processing update 8. AstroGrid liaison update 9. AOB Minutes ------- 1. Actions from last meeting JRL has finished off the OMP database access scripts. There was a bit of confusion regarding this item. It was to check that the OMP database was mirroring all the necessary tables - this has been done. JRL said that the main problem with the mirror is when the OMP database changes its IP address. The mirror is currently down, so he is looking into it and will reflect on the problem. JRL has not added an MSB quality flag to the FITS header - ongoing <<<< MJI continues to investigate the merged classification statistics - ongoing <<<< WJS was not present so it is unclear whether visiting RAL during AIT3a is still a possibility - ongoing <<<< DWE had a meeting to discuss the future design of VDFS web pages. The main EGS result from the questionnaire was to go for an all Plone system. DWE MJI and EGS are currently working on a trial system. EGS has generated some stacked DXS data and catalogues. It is likely that this will be part of the Early Data Release (EDR). MJI produced colour-colour and colour-magnitude diagrams of IC4665 and M17 along with a brief howto. This can be found at http://www.ast.cam.ac.uk/~wfcam/docs/reports/sv/#science MR was asked to make this visible from the Plone system as well. <<<< MJI asked Konrad Kuijken about the ordering of OmegaCam extensions within a FITS MEF file. Their system uses a fixed order. MJI clarified on the deliverable date for the DRL. The minimum deliverable is v0.5 by September 2006. However, it should be noted that this is relative to the date of commissioning. PSB have not come up with a suitable strategy for jittering VISTA JRL twilight flats. This is not a priority at the moment - ongoing <<<< MJI DWE released the latest PSF report http://www.ast.cam.ac.uk/~wfcam/docs/reports/psf3/ JRL has not finished the mods to AAT archive. He has asked if there is any more AAT data that needs ingesting, but has not had a reply - ongoing <<<< JRL has not asked if there is any more UKIRT data that needs ingesting - ongoing <<<< EGS have sorted out an automatic and reliable reboot for apm14 and apm15. MR MR said that PSB knows what the root password is for both machines. MR also raised the point that the current Plone system is accessed directly via their machine names, so it would not be easy to switch over in the event of a hardware failure. DWE said that this would be sorted out with the new Plone system. 2. Comments on WFAU minutes PSB noted that the UKLight options are still open and will be interested in the outcome of investigations. RGM noted that the transfers to Edinburgh are now 2 months behind and that the ESO transfers have overtaken this. STH asked what data was being released as part of the EDR. (This was discussed later - item 4) 3. VDFS VISTA update Telescope status - MJI said that the progress was very promising and that WJS gave a report at the VISTA Science Surveys meeting. VIRCAM - PSB said that Gavin Dalton had reported that they had traced the source of the stray light entering the camera. The AIT3a tests are currently in progress. PSB was uncertain what the plans for the next few months were. VDFS Documentation - PSB said that he had identified 9 documents that were pending in that they had not been signed or released yet. He announced that he was making it his personal campaign (PPC) to get these documents sorted out. He said that he would also like these documents freely available on the web rather than sitting on our internal system. DRL status - It was pointed out that v1.2 of the above documents are those that are regarded as the final FDR documents. It was also noted that as JRL advances the DRL, the appropriate documentation is being updated. JMS wants Sandra Castro to write a document on her views regarding the VIRCAM pipeline and then have a telecon to discuss this. There were a number of opinions regarding this. ETC - This new CASU requirement occupied most of PSB's (& some of MJI's) time during January. It was functional by 15 January. The main difficulty was in obtaining some of the data. Much of it had to be converted into a suitable from and some was not available. Although help pages were generated, this was handicapped by the ETC document not being available to users even though the ETC was. Workstations at Paranal - Although there has been some movement on this front, it is still not clear if more than one will be available. VDFS UK Review - MJI distributed the Terms of Reference for this review, however it doesn't say what documents are needed. It was noted that the UK User Requirements Document has not been signed off and it was now a bit out of date. A likely date for the review would be the end of April. It was not known who would be on the review panel. These issues should be raised at the next VDMT meeting. <<<< It was noted that we only have 3 FTE working on VDFS at the moment. VISTA Surveys - There have been a number of meetings recently: VISTA Hemispheres (Cambridge - PSB, MJI and STH went to it), GPS (Hatfield) and SuperDXS (will be held tomorrow at Oxford). Many of these surveys will be similar to the VST surveys. VISTA operations - It was unclear if we were only going to process the UK led surveys. 4. WFCAM processing MR said that the processing was progressing well. Processing was occurring on data up to 10 January. We are waiting for a tape with 5 more nights of data. Data up to 31 December has been processed, up to 20 October has been transfered to Edinburgh and up to 25 October has been transfered to ESO. JRL reported on the improved sky algorithm. He said that he would wait until the current batch of data has been processed and then test the algorithm on some real data (about end of January). DWE reported on persistence. A quick look indicates that any pixel that gets saturated will have a persistence effect. This can be up to about 100 counts and decays over about 60 seconds. It is unclear if there is a filter dependency. The difficulty is in identifying saturated pixels since they can have any value - unlike with a CCD. MJI said that using the catalogue data might help here. It is necessary to decide what to do about persistence before the reprocessing starts (in February). DWE was asked to provide some further results. <<<< There was some discussion regarding how much of the data was checked by eye. MJI said that a small fraction was looked at and randomly selected. JRL said that if there was a problem with a frame, 90% were caused by problems with the sky determination. RGM said that for the EDR UKIDSS data from 05A will be released. This will not include the SV data. The exception from this is that the 05B September data for DXS and UDS would also be released. Additionally there are some data quality checks further restricting this selection. The EDR paper was also discussed? DWE reported on the PSF work. The split PSF measurer is working well and a baseline PSF fitting programme using this has been written. This latter programme is not in an ideal condition and more work is needed on this. MJI said that the BBC had visited WFCAM on the 12 January and some frames for publicity had been taken. This data will be on the next batch of tapes to arrive. JRL said that he would be processing this data (needed by 2nd February). <<<< There was some discussion about if there was a need for a JAC telecon. Since there were no items to bring up, it was decided that there wasn't a need. RGM(?) asked if there was any news from the UKIRT board meeting. MJI said that he would try and find out. <<<< STH said that he, Paul Hewett, PSB and JPE had a meeting with Martin Cohen to discuss absolute photometric calibration. He has a network of absolute standards (down to 14 magnitude) which could be used for the Y and Z calibration. A misunderstanding regarding the use of 2MASS for WFCAM calibration was clarified and he now agrees that the calibration can be done to the 1-2% level. 5. Upcoming VDMT meeting and this quarter WPs STH said that an extra work package has been added to account for the extra work on the VISTA ETC. He reminded us that he wanted the monthly reports (covering December and January) in by the end of today so that MJI could review the results on Friday. <<<< The work for the UK Review will come in during this quarter, so some clarification regarding the review will be necessary from the VDMT meeting. 6. Data archives update The INT data has now been all loaded by RWA. He and EGS are sorting out the organizational changes that are required. The data was ordered by tape number, but it is now ordered by date (?). No news regarding the rest of the archives. 7. Optical processing update Most of the optical processing that is being carried out by the group at the moment is on IPHAS data. STH said that we will be processing some 200Gb of INT data for an Oxford group. This data was taken between July 2004 and January 2006. RGM said that he and a student are using DR4 and DR5 SDSS data to calibrate INT data taken under non-photometric conditions and deriving new (?) colour equations. 8. AstroGrid liaison update RGM said that the EU have awarded money for an extra person to work on AstroGrid in Cambridge for 2 years. 9. AOB NAM - This will be at Leicester (Monday 3 to Friday 7 April) and it was felt that a CASU presence was required if only to show the progress of the reductions. RGM said that he would look further into this. <<<< ADASS - JRL said that the venue has been booked for the week starting 23 September 2007. Di Sword will be helping organize the conference dinner and the reception. Science with WFCAM data was discussed. MJI was asked to find out from Andy Lawrence (CC Andy Adamson) about data rights on the processed data. <<<< Continuing Actions ------------------ JRL provide a script to add JAC MSB quality flag to the FITS headers MJI finish investigating merged classification statistics WJS advance the proposed CASU visit to RAL, in particular check that having it during AIT3a is OK PSB suggest a suitable strategy for jittering VISTA twilight flats JRL MJI JRL finish mods to AAT archive and sort out ingesting more AAT data if more needs ingesting JRL ask if there is any more UKIRT data that needs ingesting New Actions ----------- MR make the science plots of IC4665 and M17 visible from the Plone system STH raise various UK Review issues at VDMT meeting DWE provide some further results on persistence JRL process the publicity frames for the BBC (by 2 February) MJI acquire news from the UKIRT board meeting ALL send in 2 months' worth of reports to STH by end of today (26th) RGM investigate possible CASU presence at NAM 2006 MJI ask Andy Lawrence (CC Andy Adamson) about data rights on the processed data